OpenStudy (anonymous):

write the equation of a line that includes the point (1,5) and has a slope of 3 in slope intercept form

OpenStudy (texaschic101):

y = mx + b slope(m) = 3 (1,5)...x = 1 and y = 5 now we sub....and since we already know slope(m) and our points (x,y)...we need to find b, the y intercept. y = mx + b 5 = 3(1) + b 5 = 3 + b 5 - 3 = b 2 = b so your equation is : y = 3x + 2